If x=(5)^1/3+2,then value of x^3-6x^2+12x-10 is (a) 1  (b) -2. (c) -1 (d) 3

Grade:10

2 Answers

Aditya Gupta
5 years ago
we see that x-2=(5)^1/3
cubing both sides and rearranging we get
x^3-6x^2+12x=13
so x^3-6x^2+12x-10= 3
option d 3 is correct
vibha
5 years ago
sol. in question given